Which is greater in each of the following: -3/4, 2/-3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The LCM of the denominators 4 and 3 is 12
∴(-3/4) = [(-3 × 3)/ (4 × 3)]= (-9/12)
And (-2/3) = [(-2 × 4)/ (3 × 4)] = (-8/12)
Now,
-9 < -8
So, (-9/12) < (- 8/12)
∴-3/4 < 2/-3
Hence, 2/-3 is greater

​​Thanks